# Unlocking the Power of Mixture-of-Experts for Task-Aware Time Series Analytics

### v1

@@ Appendix E Related Works
Time series forecasting (TSF) involves predicting future observations by leveraging historical data. Research has indicated that features derived through learning algorithms can potentially outperform human-engineered features, as demonstrated in multiple studies (Qiu et al., 2025b; a; Liu et al., 2025a; Lu et al., 2023; 2024a; 2024b; Li et al., 2025a; Yang et al., 2024; Zhou et al., 2025; Li et al., 2025b; Wang et al., 2025; Wu et al., 2024c; 2025a; a; Qiu et al., 2025c; Wu et al., 2025b; Qiu et al., 2025e; Dai et al., 2024c; Liu et al., 2025b; Sun et al., 2025b; a; Niu et al., 2025). By capitalizing on the representation learning capabilities of deep neural networks (DNNs), numerous deep-learning approaches have emerged.
Methods such as TimesNet (Wu et al., 2023) and SegRNN (Lin et al., 2023) treat time series as sequences of vectors. They utilize convolutional neural networks (CNNs) or recurrent neural networks (RNNs) to capture temporal dependencies. Moreover, Transformer-based architectures, including Informer (Zhou et al., 2021), TimeBridge (Liu et al., 2024b), PDF (Dai et al., 2024b), Triformer (Cirstea et al., 2022), and PatchTST (Nie et al., 2023), are capable of more accurately discerning complex relationships among time points. This has led to a significant enhancement in forecasting performance.
MLP-based methods, such as DUET (Qiu et al., 2025d), CycleNet (Lin et al., 2024b), NLinear (Zeng et al., 2023), and DLinear (Zeng et al., 2023), adopt relatively simpler architectures with fewer parameters. Despite this, they still manage to achieve highly competitive levels of forecasting accuracy.

### v2

@@ Appendix E Related Works
Time Series Analysis holds paramount importance across diverse fields, including the economy (Qiu et al., 2025b; a; Liu et al., 2025a), transportation (Wu et al., 2024c; a), health (Lu et al., 2023; 2024a; 2024b), weather (Li et al., 2025a; Yang et al., 2024; Zhou et al., 2025), and energy (Sun et al., 2025a). It encompasses multiple key tasks, such as forecasting (Li et al., 2025b; Dai et al., 2024c), anomaly detection (Wang et al., 2025; Qiu et al., 2025c; Wu et al., 2025b), classification (Liu et al., 2024d), imputation (Wu et al., 2023), and others (Qiu et al., 2025e; Wu et al., 2025a). Among these, Time Series Forecasting is the most extensively applied in real-world scenarios.
Time series forecasting (TSF) entails predicting future observations by exploiting historical data. Research findings suggest that features derived from learning algorithms have the potential to outperform human-engineered features, as evidenced by numerous studies (Liu et al., 2025b; Sun et al., 2025b; Niu et al., 2025). Leveraging the representation learning capabilities of deep neural networks (DNNs), a plethora of deep-learning approaches have emerged.
Methods like TimesNet (Wu et al., 2023) and SegRNN (Lin et al., 2023) treat time series as sequences of vectors. They utilize convolutional neural networks (CNNs) or recurrent neural networks (RNNs) to capture temporal dependencies. Furthermore, Transformer-based architectures, including Informer (Zhou et al., 2021), TimeBridge (Liu et al., 2024b), PDF (Dai et al., 2024b), Triformer (Cirstea et al., 2022), and PatchTST (Nie et al., 2023), are adept at discerning complex relationships among time points more accurately. This has significantly enhanced forecasting performance.
MLP-based methods, such as DUET (Qiu et al., 2025d), CycleNet (Lin et al., 2024b), NLinear (Zeng et al., 2023), and DLinear (Zeng et al., 2023), adopt relatively simpler architectures with fewer parameters. Nevertheless, they can still attain highly competitive levels of forecasting accuracy.